The perimeter of a rectangle is 524 in. The length is 42in. more than the width. What is the length?
Answer by user_dude2008(1862)

L=42+W

2L+2W=P
2(42+W)+2W=524
84+2W+2W=524
2W+2W=524-84
4W=440
W=440/4
W=110 <------------- Width is 110 in
L=42+W=42+110=152 <------------- Length is 152 in